# DeepEye-SQL: A Software-Engineering-Inspired Text-to-SQL Framework

### v2

@@ Abstract.
Large language models (LLMs) have advanced Text-to-SQL, yet existing solutions still fall short of system-level reliability. The limitation is not merely in individual modules – e.g., schema linking, reasoning, and verification – but more critically in the lack of structured orchestration that enforces correctness across the entire workflow. This gap motivates a paradigm shift: treating Text-to-SQL not as free-form language generation but as a software-engineering problem that demands structured, verifiable orchestration. We present DeepEye-SQL, a software-engineering-inspired framework that reframes Text-to-SQL as the development of a small software program, executed through a verifiable process guided by the Software Development Life Cycle (SDLC). DeepEye-SQL integrates four synergistic stages: it grounds ambiguous user intent through semantic value retrieval and robust schema linking; enhances fault tolerance with N-version SQL generation using diverse reasoning paradigms; ensures deterministic verification via a tool-chain of unit tests and targeted LLM-guided revision; and introduces confidence-aware selection that clusters execution results to estimate confidence and then takes a high-confidence shortcut or runs unbalanced pairwise adjudication in low-confidence cases, yielding a calibrated, quality-gated output. This SDLC-aligned workflow transforms ad hoc query generation into a disciplined engineering process. Using $\sim$30B open-source LLMs without any fine-tuning, DeepEye-SQL achieves 73.5% execution accuracy on BIRD-Dev and 89.8% on Spider-Test, outperforming state-of-the-art solutions. This highlights that principled orchestration, rather than LLM scaling alone, is key to achieving system-level reliability in Text-to-SQL.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Figure 1. Key Idea of DeepEye-SQL
Text-to-SQL is a task that converts natural-language questions into SQL queries over a database (Luo et al., 2025b; Liu et al., 2025a; Luo et al., 2018, 2022, 2021, 2025a; Baek et al., 2025). Large language models (LLMs) (Wu et al., 2025; Shuai et al., 2025; Zhu et al., 2024; Zhang et al., 2025d) have substantially advanced Text-to-SQL, achieving strong results on benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2024a). For example, Alpha-SQL (Li et al., 2025b) leverages dynamic multi-step reasoning, while XiYan-SQL (Liu et al., 2025c) improves SQL generation and multi-candidate SQL selection through task-specific fine-tuning.
Despite these advances, state-of-the-art performance on the BIRD dataset remains around 70% execution accuracy (Li et al., 2024a), and reliability further degrades in real-world deployments (Li et al., 2024b; Chung et al., 2025). This observation indicates that recent advances, while promising, have yet to translate into consistent system-level reliability.
The key limitation lies not in the optimization of individual modules – such as schema linking, reasoning, or post-hoc verification – but in the lack of coherent orchestration that enforces correctness across the entire workflow (Liu et al., 2025c; Li et al., 2025b; Pourreza and Rafiei, 2023; Xie et al., 2025; Gao et al., 2024; Talaei et al., 2024; Pourreza et al., 2025). As a result, current Text-to-SQL solutions struggle to: (i) define what should be built, which requires precisely determining the semantic scope of the user question and grounding it to the relevant database entities through comprehensive schema linking and value retrieval; (ii) implement the solution (i.e., SQL generation), which involves generating executable SQL queries that faithfully capture the inferred semantics while maintaining diversity and completeness across complex reasoning paths; (iii) verify its correctness, which requires systematically validating the structural, logical, and semantic correctness of the generated SQL through interpretable checks (Liu et al., 2025b); and (iv) release the generated SQL, which requires quantifying confidence through multi-source evidence and establishing measurable acceptance criteria for determining whether a generated SQL query is reliable enough for output.
@@
First, Text-to-SQL solutions must infer user intent from ambiguous natural language and partially observed database schemas. We term this challenge ambiguous requirement inference (challenge C1). In the implementation stage, current methods rely on a single reasoning path driven by one model (Pourreza and Rafiei, 2023) and one prompt configuration (Li et al., 2024b), resulting in insufficient fault tolerance (C2). In the verification and validation stage, existing methods depend on probabilistic signals such as self-consistency (Li et al., 2025b) or partial execution feedback (Zhang et al., 2025a) rather than deterministic oracles to assess generated SQL. We denote this challenge as unreliable verification and validation (C3). In the release stage, current methods lack calibrated confidence estimation and measurable acceptance criteria for evaluating when a generated SQL query is reliable enough for output (C4).
Our Methodology: Software-Engineering-Inspired Framework. To systematically address these challenges, we propose DeepEye-SQL, which reframes Text-to-SQL as a verifiable SDLC-style workflow with four stages (Figure 1). In requirements analysis, we propose Semantic Value Retrieval and Robust Schema Linking, i.e., combining direct, reversed, and value-based linking with relational closure, to build a complete, database-grounded specification, addressing ambiguous requirement inference (addressing the challenge C1). In implementation, we adopt N-version programming (Chen and Avizienis, 1978) with three independent generators (skeleton, in-context learning, and divide-and-conquer) executed in parallel to diversify reasoning under a fixed budget, providing fault tolerance (addressing C2). For verification and validation, we replace probabilistic self-judgment with a tool-chain of checkers (syntax, clause/time, JOIN, NULL/result) that trigger targeted LLM repair, ensuring verifiable correctness (addressing C3). Finally, in release, we introduce confidence-aware selection that clusters execution results to estimate confidence and then takes a high-confidence shortcut or runs unbalanced pairwise adjudication in low-confidence cases, yielding a calibrated, quality-gated output (addressing C4). As shown in Figure 2, this design enables DeepEye-SQL to integrate with diverse LLMs and achieve notable accuracy improvements without any fine-tuning.
@@
(1) A Novel SE-Inspired and Training-free Framework. We propose DeepEye-SQL, which reframes Text-to-SQL as an SDLC-guided, four-stage engineering workflow. The framework is training-free and plug-and-play, improving system-level reliability across diverse LLMs without any fine-tuning.
(2) A Suite of Principled Techniques for Reliability. We introduce a set of techniques, each grounded in a specific software engineering principle, to ensure end-to-end reliability. This includes: (i) a fault-tolerant schema linking mechanism that combines multiple strategies to guarantee specification completeness; (ii) a deterministic, tool-based revision process that acts as a unit testing suite for reliable debugging; and (iii) an adaptive, confidence-aware selection mechanism that serves as a quality gate for the final output.
@@
(3) Extensive Experiments demonstrate that DeepEye-SQL achieves state-of-the-art performance on challenging benchmarks. With the Qwen3-Coder-30B-A3B model, DeepEye-SQL attains an execution accuracy of 73.5% on BIRD-Dev and 89.8% on Spider-Test without any model fine-tuning.
Figure 2. DeepEye-SQL, a plug-and-play Text-to-SQL framework, consistently surpasses prior SOTA methods using $\sim$30B open-source LLMs without any task-specific fine-tuning.
2. Problem Definition and Preliminary
2.1. Problem Definition
Text-to-SQL Task. The primary objective of the Text-to-SQL task is to automatically translate a user’s question, expressed in natural language, into a corresponding SQL query. Formally, given a database schema $\mathcal{D}$ and a natural language question $\mathcal{Q}$, the task is to generate a SQL query $\mathcal{S}$ that correctly retrieves the answer to $\mathcal{Q}$ from the database instance defined by $\mathcal{D}$. The schema $\mathcal{D}$ is defined as a set of tables $\{T_{1},T_{2},\dots,T_{m}\}$, where each table $T_{i}$ consists of a set of columns $\{C_{i1},C_{i2},\dots,C_{ik}\}$. The goal is to find a mapping function $f$ such that:
$\mathcal{S}=f(\mathcal{Q},\mathcal{D})$
Let $\llbracket\mathcal{S}\rrbracket_{\mathcal{D}}$ denote the result of executing query $\mathcal{S}$ on the database. The generated query $\mathcal{S}$ is considered correct if $\llbracket\mathcal{S}\rrbracket_{\mathcal{D}}$ accurately answers the user’s intent expressed in $\mathcal{Q}$. The core challenge lies in bridging the semantic gap between the unstructured natural language $\mathcal{Q}$ and the structured query language representation $\mathcal{S}$ within the context of the given schema $\mathcal{D}$.
2.2. Text-to-SQL Solutions
The task of translating natural language into executable SQL queries has evolved significantly (Liu et al., 2025a). The recent emergence of Large Language Models (LLMs) has marked a new paradigm. Current research on applying LLMs to Text-to-SQL is largely divided into two categories: fine-tuning and prompting-based methods. Fine-tuning methods adapt open-source models like Code Llama or Qwen for the Text-to-SQL task by training them on large corpora of question-SQL pairs (Li et al., 2024c; Yang et al., 2024; Li et al., 2025a; Sheng et al., 2025; Liu et al., 2025c). This approach can yield highly efficient and specialized models but may exhibit limited generalization to out-of-domain or highly complex scenarios. In contrast, prompting-based methods leverage the powerful in-context learning and reasoning capabilities of very large, often proprietary, models like GPT-4o (Hurst et al., 2024) and Gemini (Anil et al., 2023) without requiring model training. To manage the complexity of the task, these methods typically decompose the process into a multi-stage pipeline, including sub-tasks like schema linking, SQL generation, and refinement (Pourreza and Rafiei, 2023; Li et al., 2024b; Pourreza et al., 2025; Li et al., 2025b; Ma et al., 2024; Zhu et al., 2025). While powerful, these frameworks often rely on a single generation path and the fallible self-correction abilities of the LLM, which can compromise robustness. Our work, DeepEye-SQL, belongs to the prompting-based category but differentiates itself by explicitly adopting a systematic framework inspired by software engineering.
2.3. Software Engineering
The challenge of building robust and reliable LLM-based systems often mirrors the complexities of traditional software development. In response, software engineering has established a set of core principles to manage complexity and ensure product quality. The Software Development Life Cycle (SDLC) (Ruparelia, 2010) provides a foundational, systematic process for software creation, typically involving phases such as requirements analysis, implementation, testing, and deployment. To deconstruct existing systems and inform new designs, practitioners often employ Reverse Engineering (Rekoff, 1985), a process of analyzing a finished product to deduce its underlying specifications. To enhance system reliability, fault tolerance techniques are critical. A notable example is N-Version Programming (Chen and Avizienis, 1978), where multiple, independently developed versions of a component are executed in parallel, and their results are adjudicated to mask faults and increase the likelihood of a correct outcome. For quality assurance, Unit Testing (Runeson, 2006) is a fundamental practice. It involves the granular testing of individual software components or “units” in isolation to verify that each part functions correctly according to its design specifications. This bottom-up approach is crucial for identifying and rectifying defects early in the development process. Finally, to govern the release process, a Quality Gate (Schnell and Reinhart, 2016) acts as a final checkpoint, enforcing a set of predefined criteria to determine whether a software artifact meets the required quality standard for deployment. While these principles are cornerstones of traditional software engineering, their systematic application to the construction of LLM-based pipelines is an emerging area. Our work is directly inspired by these paradigms, leveraging them to bring a more structured, verifiable, and robust engineering discipline to the Text-to-SQL problem.
@@ 3. DeepEye-SQL Overview
3.1. Overall Architecture
Figure 3. DeepEye-SQL Overview.:
@@ 3.1. Overall Architecture
Phase-1: Intent Scoping and Semantic Grounding. Mirroring the initial phase of any engineering project—answering “What should be built?”—this stage is dedicated to accurately interpreting user intent and defining the problem’s scope. It employs Semantic Value Retrieval to ground the query in the database’s actual data and Robust Schema Linking to identify necessary tables and columns. This linking module uses a fault-tolerant hybrid strategy, combining Direct, Reversed (inspired by reverse engineering (Rekoff, 1985)), and Value-based methods avoid the “single point of failure” problem (Ulbrich et al., 2012), ensuring a comprehensive and fault-tolerant specification.
@@
Phase-2: N-version Programming for SQL Generation. Analogous to the implementation phase, this stage generates SQL queries. To enhance robustness, we employ a strategy inspired by N-version programming (Chen and Avizienis, 1978), a fault-tolerance technique where multiple, independent implementations are created for the same problem. Our framework instantiates this by producing a diverse set of SQL candidates in parallel from three distinct generators: a Skeleton-based, an ICL-based, and a Divide-and-Conquer-based generator. It is crucial to distinguish this approach from test-time scaling techniques like self-consistency (Wang et al., 2023), which generate diversity by sampling multiple outputs from a single generator. In contrast, our method achieves a more principled and profound diversity, akin to true N-version programming, as each generator employs a fundamentally different reasoning process (e.g., SQL skeleton planning, example-based reasoning, recursive decomposition). This engineered diversity ensures a broader exploration of the solution space, significantly increasing the probability of producing at least one correct query, especially for complex scenarios where a single reasoning path might fail.
@@
Phase-3: SQL Unit Testing and Revision via Tool-Chain. This phase embodies the critical software engineering principle of Unit Testing (Runeson, 2006). Its purpose is to systematically verify the correctness of each generated SQL candidate and revise any found defects. To overcome the known unreliability of LLM self-correction (Zhang et al., 2025c), our framework externalizes this process, emulating a rigorous, automated testing loop. Each SQL candidate is passed through a Tool-Chain of Checkers—a suite of specialized, deterministic tools where each checker acts as a test case for a specific unit of functionality (e.g., syntax, JOIN correctness). If a flaw is detected, the tool provides an explicit and actionable directive to the LLM for a targeted revision, mirroring a formal bug report and fix cycle.
@@
Phase-4: Confidence-aware SQL Selection. The final phase corresponds to the release stage, governed by a Quality Gate. Instead of simply choosing the most common answer, this stage arbitrates which candidate is reliable enough to be “released”. Our Confidence-aware SQL Selection mechanism performs this task. It first calculates an execution-based confidence score for the top candidates. This score is then evaluated against a predefined confidence threshold to validate reliability. This validation determines the execution path: high-confidence queries (those exceeding the threshold) pass the quality gate directly for efficiency, while low-confidence queries trigger a more nuanced, unbalanced LLM voting process to ensure the most reliable query is ultimately selected.

### v3

@@ Abstract.
Large language models (LLMs) have advanced Text-to-SQL, yet existing solutions still fall short of system-level reliability. The limitation is not merely in individual modules – e.g., schema linking, reasoning, and verification – but more critically in the lack of structured orchestration that enforces correctness across the entire workflow. This gap motivates a paradigm shift: treating Text-to-SQL not as free-form language generation but as a software-engineering problem that demands structured, verifiable orchestration. We present DeepEye-SQL, a software-engineering-inspired framework that reframes Text-to-SQL as the development of a small software program, executed through a verifiable process guided by the Software Development Life Cycle (SDLC). DeepEye-SQL integrates four synergistic stages: it grounds user intent through robust schema linking, enforcing relational closure; enhances fault tolerance with N-version SQL generation; ensures deterministic verification via a “Syntax-Logic-Quality” tool-chain that intercepts errors pre-execution; and introduces confidence-aware selection that leverages execution-guided adjudication to resolve ambiguity beyond simple majority voting. Leveraging open-source MoE LLMs ($\sim$30B total, $\sim$3B activated parameters) without any fine-tuning, DeepEye-SQL achieves 73.5% execution accuracy on BIRD-Dev, 75.07% on the official BIRD-Test leaderboard, and 89.8% on Spider-Test, outperforming state-of-the-art solutions that rely on larger models or extensive training. This highlights that principled orchestration, rather than LLM scaling alone, is key to achieving system-level reliability in Text-to-SQL.
Text-to-SQL, Databases, Large language models
copyright: nonejournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 3 (SIGMOD)article: 158publicationmonth: 6doi: 10.1145/3802035ccs: Information systems Structured Query Language
@@ 1. Introduction
Figure 1. Key Idea of DeepEye-SQL.
Table 1. Comparison of DeepEye-SQL with representative SOTA Text-to-SQL methods.
Method Core Paradigm Verification Mechanism SQL Selection Strategy Efficiency & Cost Profile
CHESS (Talaei et al., 2024) Linear Inference Workflow LLM Self-Refinement Heuristic Majority Voting High API Cost
(Single-path, vulnerable to error propagation) (Probabilistic, hallucination-prone) (Standard Self-Consistency) (Gemini 1.5 Pro, Proprietary)
Alpha-SQL (Li et al., 2025a) Search-Centric (MCTS) Basic Syntax Check Heuristic Majority Voting High Inference Latency
(Exploration via Monte Carlo Search) (Ensures executability only) (Relies on rollout counts) (MCTS requires massive inferences)
OmniSQL (Li et al., 2025b) Data-Centric SFT Basic Syntax Check Heuristic Majority Voting High Training Cost
(Synthetic Data Augmentation) (Ensures executability only) (Standard Self-Consistency) (Requires Full Fine-Tuning)
[…]
Text-to-SQL is a task that converts natural-language questions into SQL queries over a database (Luo et al., 2025b; Liu et al., 2025a; Luo et al., 2018, 2022, 2021, 2025a; Baek et al., 2025; Zhu et al., 2025b, a). Large language models (LLMs) (Wu et al., 2025; Shuai et al., 2025; Zhu et al., 2024; Zhang et al., 2025a; Xie et al., 2025b, 2024) have substantially advanced Text-to-SQL, achieving strong results on benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2024c). For example, Alpha-SQL (Li et al., 2025a) leverages dynamic multi-step reasoning, while XiYan-SQL (Liu et al., 2025c) improves SQL generation and multi-candidate SQL selection through task-specific fine-tuning.
Despite these advances, state-of-the-art performance on the BIRD dataset remains around 70% execution accuracy (Li et al., 2024c), and reliability further degrades in real-world deployments (Li et al., 2024a; Chung et al., 2025; Zhang et al., 2021). This observation indicates that recent advances, while promising, have yet to translate into consistent system-level reliability.
The key limitation lies not in the optimization of individual modules – such as schema linking, reasoning, or post-hoc verification – but in the lack of coherent orchestration that enforces correctness across the entire workflow (Liu et al., 2025c; Li et al., 2025a; Pourreza and Rafiei, 2023; Xie et al., 2025a; Gao et al., 2024; Talaei et al., 2024; Pourreza et al., 2025). As a result, current Text-to-SQL solutions struggle to: (i) define what should be built, which requires precisely determining the semantic scope of the user question and grounding it to the relevant database entities through comprehensive schema linking and value retrieval; (ii) implement the solution (i.e., SQL generation), which involves generating executable SQL queries that faithfully capture the inferred semantics while maintaining diversity and completeness across complex reasoning paths; (iii) verify its correctness, which requires systematically validating the structural, logical, and semantic correctness of the generated SQL through interpretable checks (Liu et al., 2025b); and (iv) release the generated SQL, which requires quantifying confidence through multi-source evidence and establishing measurable acceptance criteria for determining whether a generated SQL query is reliable enough for output.
@@
First, Text-to-SQL solutions must infer user intent from ambiguous natural language and partially observed database schemas. We term this challenge ambiguous requirement inference (challenge C1). In the implementation stage, current methods rely on a single reasoning path driven by one model (Pourreza and Rafiei, 2023) and one prompt configuration (Li et al., 2024a), resulting in insufficient fault tolerance (C2). In the verification and validation stage, existing methods depend on probabilistic signals such as self-consistency (Li et al., 2025a) or partial execution feedback (Zhang et al., 2025d) rather than deterministic oracles to assess generated SQL. We denote this challenge as unreliable verification and validation (C3). In the release stage, current methods lack calibrated confidence estimation and measurable acceptance criteria for evaluating when a generated SQL query is reliable enough for output (C4).
Our Methodology. To systematically address these challenges, we propose DeepEye-SQL, which reframes Text-to-SQL as a verifiable SDLC-style workflow with four stages (Figure 1). In requirements analysis, we propose Semantic Value Retrieval and a fault-tolerant Robust Schema Linking strategy to build a complete, database-grounded specification (addressing C1). In implementation, we adopt N-version programming (Chen and Avizienis, 1978) by executing multiple independent generators with diverse reasoning paradigms in parallel to diversify reasoning under a fixed budget, providing fault tolerance (addressing C2). For verification and validation, we replace probabilistic self-judgment with a suite of specialized, deterministic checkers that trigger targeted LLM repair, ensuring verifiable correctness (addressing C3). Finally, in release, we introduce Confidence-aware SQL Selection that dynamically adapts the review process based on execution consensus, yielding a calibrated output (addressing C4). As shown in Figure 2, this design enables DeepEye-SQL to integrate with diverse LLMs and achieve notable accuracy improvements without any fine-tuning. To explicitly distinguish our contribution from existing approaches, Table 1 contrasts DeepEye-SQL with representative state-of-the-art methods. Unlike linear workflows (e.g., CHESS (Talaei et al., 2024)) or search-centric methods (e.g., Alpha-SQL (Li et al., 2025a)) that rely heavily on single-path reasoning or probabilistic self-correction, DeepEye-SQL enforces a deterministic SE-Lifecycle paradigm. We also incorporate a fault-tolerant design via N-Version Programming to mitigate single-point reasoning failures. Furthermore, by replacing stochastic checks with a rigorous Unit Testing Tool-Chain and introducing Confidence-Aware Selection, we systematically address the inherent unreliability of LLM self-reflection.
Figure 2. As a plug-and-play framework, DeepEye-SQL consistently outperforms SOTA baselines using $\sim$30B LLMs.
@@
(1) A Novel SE-Inspired Framework. We propose DeepEye-SQL, which reframes Text-to-SQL as a verifiable SDLC workflow. Unlike prior frameworks constrained by fragile, linear inference chains where errors propagate unchecked, DeepEye-SQL enforces a robust engineering lifecycle (Requirement $\to$ Implementation $\to$ Testing $\to$ Release) to guarantee system-level reliability.
(2) N-Version Programming for Diversity. Departing from standard self-consistency methods that rely on stochastic sampling (same prompt, high temperature), we introduce an N-Version Programming paradigm. By orchestrating independent generators with distinct reasoning strategies, we achieve true algorithmic fault tolerance and broader coverage of complex queries.
(3) Deterministic Verification via Tool-Chain. To address the confirmation bias inherent in LLM self-correction, we propose a novel SQL Unit Testing Tool-Chain. This module replaces probabilistic LLM self-reflection with deterministic, external execution checks (Syntax, Logic, and Data Quality), providing precise feedback that grounds the revision process.
@@
(4) Extensive Experiments demonstrate that DeepEye-SQL achieves state-of-the-art performance on challenging benchmarks. Leveraging open-source MoE LLMs with only $\sim$3B activated parameters, DeepEye-SQL attains an execution accuracy of 73.5% on BIRD-Dev, 75.07% on the held-out BIRD-Test set, and 89.8% on Spider-Test, outperforming larger or fine-tuned baselines.
@@ 2. DeepEye-SQL Overview
Figure 3. DeepEye-SQL Overview.
@@
Phase-1: Intent Scoping and Semantic Grounding. Mirroring the initial phase of any engineering project—answering “What should be built?”—this stage is dedicated to accurately interpreting user intent and defining the problem’s scope. It employs Semantic Value Retrieval to ground the query in the database’s actual data and Robust Schema Linking to identify necessary tables and columns. This linking module uses a fault-tolerant hybrid strategy that integrates complementary linking paradigms to avoid the “single point of failure” problem (Ulbrich et al., 2012), ensuring a comprehensive and fault-tolerant specification.
@@
Phase-2: N-version Programming for SQL Generation. Analogous to the implementation phase, this stage generates SQL queries. To enhance robustness, we employ a strategy inspired by N-version programming (Chen and Avizienis, 1978), a fault-tolerance technique where multiple, independent implementations are created for the same problem. Our framework instantiates this by producing a diverse set of SQL candidates in parallel from three distinct generators, each utilizing a unique reasoning paradigm. It is crucial to distinguish this approach from test-time scaling techniques like self-consistency (Wang et al., 2023), which generate diversity by sampling multiple outputs from a single generator. In contrast, our method achieves a more principled and profound diversity, akin to true N-version programming, as each generator employs a fundamentally different reasoning process. This engineered diversity ensures a broader exploration of the solution space, significantly increasing the probability of producing at least one correct query, especially for complex scenarios where a single reasoning path might fail.
@@
Phase-3: SQL Unit Testing and Revision via Tool-Chain. This phase embodies the critical software engineering principle of Unit Testing (Runeson, 2006). Its purpose is to systematically verify the correctness of each generated SQL candidate and revise any found defects. To overcome the known unreliability of LLM self-correction (Zhang et al., 2025b), our framework externalizes this process, emulating a rigorous, automated testing loop. Each SQL candidate is passed through a Tool-Chain of Checkers—a suite of specialized, deterministic tools where each checker acts as a test case for a specific unit of functionality. These checkers systematically validate structural and semantic constraints, and if a flaw is detected, provide an explicit and actionable directive to the LLM for a targeted revision, mirroring a formal bug report and fix cycle.
@@
Phase-4: Confidence-aware SQL Selection. The final phase corresponds to the release stage, governed by a Quality Gate. Instead of simply choosing the most common answer, this stage arbitrates which candidate is reliable enough to be “released”. Our Confidence-aware SQL Selection mechanism performs this task. It estimates the reliability of candidates based on execution consensus and evaluates this against a predefined acceptance criterion. This validation acts as a gateway, determining whether to release the result directly or trigger a deeper, comparative review to ensure the most reliable query is selected.